# Sharpness of the phase transition for constrained-degree percolation

### v1

@@ 1 Introduction
Percolation models have been extensively studied for decades. Since their introduction in [Broadbent57], they have become central objects of research in probability and mathematical physics, with numerous variants of the original model proposed over time. Several models (see e.g. [Jahnel23, Lichev24, Peled19, Garet18, Pete08, Grimmett10, Grimmett17, Holroyd21, Li20]) purposefully feature degree constraints, often making their analysis intricate. Our model of interest is the constrained degree percolation (CDP) introduced in [Teodoro14]. In CDP, each vertex may accept connections up to a fixed limit, after which no new edges can be formed, while all previously established connections remain intact. Specifically, it is a dependent continuous-time percolation process, defined by the following dynamics. At time zero, all edges are closed; each edge attempts to open at a uniformly distributed random time, and the attempt is successful if, at that moment, both of its end-vertices have degrees strictly smaller than a given parameter $\kappa$. A formal definition of the model is provided in Section 1.1. Unlike classical Bernoulli percolation, the CDP exhibits dependencies of all orders at any fixed time. Furthermore, the absence of the Fortuin–Kasteleyn–Ginibre (FKG) inequality and the lack of the finite energy property make its analysis challenging.
@@ 1.2 Background
For $\kappa=2d$, the CDP reduces to the standard Bernoulli percolation model, for which Theorem 1 is a classical result [Aizenman87, Duminil-Copin16a, Menshikov86]. We direct the reader to [Grimmett99] for some background on this model.
It is known that the CDP undergoes a nontrivial phase transition for all $d\geq 2$ and most values of $\kappa$. For example, [Delima20] establishes this non triviality for $d=2$ and $\kappa=3$, while also proving that percolation does not occur when $d\geq\kappa=2$, even at $p=1$. Further progress was made in [Hartarsky22constrainedperco], which provides quantitative upper bounds on the critical time and characterizes the phase transition for all $d\geq 3$ and most choices of $\kappa\geq 3$. Nonetheless, it remains open to determine whether the phase transition is trivial ($p_{\mathrm{c}}=1$) for $\kappa\in\{3,\dots,9\}$ for some dimensions, but it is conjectured that this is never the case. In view of this, let us emphasise that Theorem 1 applies regardless of whether the phase transition is trivial.
It is straightforward that the CDP is stochastically dominated by standard Bernoulli percolation for every fixed $p\in[0,1]$. Consequently, Theorem 1 holds trivially whenever $p$ lies below the Bernoulli percolation critical threshold, denoted by $\overline{p_{\mathrm{c}}}$. However, it is known that $p_{\mathrm{c}}>\overline{p_{\mathrm{c}}}$ (see Theorem 1 in [Delima20] for the case $d=2$ and $\kappa=3$; the argument extends for the general case), and therefore Theorem 1 remains nontrivial for all choices of $d$ and $\kappa$.
Let $\widehat{p_{\mathrm{c}}}=\sup\{p\in[0,1]:\sum_{x\in{\mathbb{Z}}^{d}}{\mathbb{P}}(0\xleftrightarrow{p}x)<\infty\}\leq p_{\mathrm{c}}$ denote the suscptibility critical threshold for the CDP. The recent work [DosSantos25] proves Theorem 1 for $p<\widehat{p_{\mathrm{c}}}$. Our result strengthens their findings: in particular, Theorem 1 implies $p_{\mathrm{c}}=\widehat{p_{\mathrm{c}}}$. This answers [DosSantos25]*Question 1. in the setting of homogeneous (vertex-independent) constraints $\kappa$.
@@ 1.3 Outline of the proof
A standard approach to proving sharpness in percolation models is the one of [Duminil-Copin19] using a randomised algorithm (see Section 2.4) and the OSSS inequality [ODonnell05] (see Section 2.3). Applying this method directly fails. On the one hand, the measure $\omega_{p}\circ{\mathbb{P}}$ is not monotonic and lacks other good properties. On the other hand, if we choose to work with the configuration space $0,1]^{\mathcal{E}}$ instead of $\{0,1\}^{\mathcal{E}}$, edges revealed in the exploration of the occurrence of the one-arm event in (1) do not witness this one-arm event on a smaller scale. For this reason, as in [Hartarsky21a], we introduce an alternative one-arm event (see Section 2.1) for which revealment is naturally expressed in terms of its probability. Namely, this event requires a path whose first part (head) is open and whose second part (tail) has decreasing values of the uniform variables on edges. This modified one-arm event satisfies a Russo formula (see Lemma 5).
Putting the above ingredients together, one reaches the main difficulty: the Russo formula features a different pivotality event ($p$-pivotality) from the influences appearing in the OSSS inequality ($U$-pivotality). In Proposition 7, we provide a transfer from one to the other. The statement is largely inspired by the essential enhancement technique of [Aizenman91]. However, the lack of finite-energy property and the infinite range of dependence make its implementation very delicate (see e.g. [Duminil-Copin18] for a very different setting where an intricate unbounded range pivotality transfer is also performed). The base idea of the proof of Proposition 7 is simple. The probability that the state of an edge depends on the uniform variables far from it decays super-exponentially, while configuration modifications usually come at an exponential cost. We can then hope to modify the configuration in the neighbourhood of the decreasing cluster of a $U$-pivotal edge in such a way that we produce a $p$-pivotal edge. Unfortunately, there exist clusters whose probability of being decreasing is only exponentially small in their volume (e.g. the fractal ternary square space-filling tree).
@@ 2.2 Transforming pivotality
Write $\tau_{n}(p)={\mathbb{P}}(E_{n}(p))$. The notion of $p$-pivotality is motivated by the following Russo formula.
@@
Any good configuration $U^{\prime}$ satisfies $\omega_{p}(U^{\prime},f)=\omega^{+}_{f}$ for all $f\in\mathcal{E}\setminus\mathcal{E}_{e}$, and $\omega_{p}(U^{\prime},f)=\mathbbm{1}_{U^{\prime}_{f}\leq p}$ for $f\in\mathcal{E}_{e}$.
@@
Let $\mathcal{E}_{+}=\{f\in\mathcal{E}_{e}:\omega^{+}_{f}=1\}$. Let $\mathcal{F}$ be a maximal subset of $\mathcal{E}_{+}$, satisfying that no good configuration $U^{\prime}$ with $\omega_{p}(U^{\prime},f)=0$ for $f\in\mathcal{E}_{+}\setminus\mathcal{F}$ belongs to $E_{n}(p)$. By Claims 9 and 10, $\mathcal{F}$ exists and $\mathcal{F}\neq\mathcal{E}_{+}$. Moreover, by the assumption of Case A, no witness of a good configuration in $E_{n}(p)$ can have a tail intersecting $\mathcal{V}_{e}$. Therefore, recalling Claim 8, within good configurations $U^{\prime}$, the event $U^{\prime}\in E_{n}(p)$ depends only on $\mathbbm{1}_{U^{\prime}_{f}\leq p}$ for $f\in\mathcal{E}_{+}$. Therefore, for any good configuration $U^{\prime}$ such that $\mathbbm{1}_{U^{\prime}_{f}\leq p}=\mathbbm{1}_{f\in\mathcal{F}}$ for all $f\in\mathcal{E}_{+}$, we have that each $f\in\mathcal{E}_{+}\setminus\mathcal{F}$ is $p$-pivotal for $U^{\prime}$ for the event $E_{n}(p)$ (recall Definition 4) and, in particular, such $f$ exist.
@@ 2.3 The OSSS inequality
An algorithm $\mathbf{A}$ determining $f$ reveals the coordinates of $x$ sequentially, with each choice depending on the values revealed thus far. The process terminates once the value of $f(x)$ is determined, independently of the unrevealed coordinates. For a formal definition of randomized algorithms, we refer to [ODonnell05]. Associated to $\mathbf{A}$ and $f$, we define the revealment and influence of coordinate $i\in I$ by
@@
where $x^{i}$ is obtained from $x$ by resampling the $i$-th coordinate independently according to $\pi$, leaving the others unchanged. The OSSS inequality [ODonnell05] states that for any Boolean function $f:X^{I}\to\{0,1\}$ and any algorithm $\mathbf{A}$ that determines $f$,
@@
This inequality was initially formulated for finite sets $X$ and $I$. Nonetheless, the result remains valid when $X$ is a general space and $I$ is countable (see [Duminil-Copin19a]*Remark 5). Applying (5) to $X=[0,1]$, $\pi$ the Lebesgue measure over $X$ and $I=\mathcal{E}$, so that $\pi^{\otimes I}={\mathbb{P}}$, and the function $f=\mathbbm{1}_{E_{n}(p)}$, and setting $\delta(\mathbf{A})=\max_{e\in\mathcal{E}}\delta_{e}(\mathbf{A})$, we obtain
@@ 2.4 The randomized algorithm approach
In this section, we apply the method of [Duminil-Copin19]*Section 3 to produce a randomized algorithm with low revealment and deduce a sharpness result for the modified one-arm event $E_{n}(p)$. Since there is little novelty here, the presentation is rather concise and we recommend referring to [Duminil-Copin19] for more details.
Lemma 13 (Differential inequality implies sharpness [Duminil-Copin19]*Lemma 3.1).

### v2

@@ 1 Introduction
Percolation models have been extensively studied for decades. Since their introduction in [4], they have become central objects of research in probability and mathematical physics, with numerous variants of the original model proposed over time. Several models (see e.g. [18, 20, 24, 11, 25, 12, 13, 17, 19]) purposefully feature degree constraints, often making their analysis intricate. Our model of interest is the constrained degree percolation (CDP) introduced in [26]. In CDP, each vertex may accept connections up to a fixed limit, after which no new edges can be formed, while all previously established connections remain intact. Specifically, it is a dependent continuous-time percolation process, defined by the following dynamics. At time zero, all edges are closed; each edge attempts to open at a uniformly distributed random time, and the attempt is successful if, at that moment, both of its end-vertices have degrees strictly smaller than a given parameter $\kappa$. A formal definition of the model is provided in Section 1.1. Unlike classical Bernoulli percolation, the CDP exhibits dependencies of all orders at any fixed time. Furthermore, the absence of the Fortuin–Kasteleyn–Ginibre (FKG) inequality and the lack of the finite energy property make its analysis challenging.
@@ 1.1 Model and result
As it is common in percolation, once sharpness is established, many questions about the subcritical phase become readily accessible. We illustrate this by the following result on the exponential decay of the volume of the cluster of the origin.
Corollary 3 (Volume exponential decay).
Consider the CDP on ${\mathbb{Z}}^{d}$, $d\geq 2$, $\kappa\in\{2,\dots,2d\}$, $p<p_{\mathrm{c}}$. Set
$C_{0}=\{x\in{\mathbb{Z}}^{d}:0\xleftrightarrow{p}x\}.$
Then there exists $\gamma_{p}>0$ such that for all $n\geq 1$,
${\mathbb{P}}(|C_{0}|\geq n+1)\leq e^{-n\gamma_{p}}.$
Corollary 3 is proved in Appendix A.
@@ 1.2 Background
For $\kappa=2d$, the CDP reduces to the standard Bernoulli percolation model, for which Theorem 1 is a classical result [1, 10, 22]. We direct the reader to [14] for some background on this model.
It is known that the CDP undergoes a nontrivial phase transition for all $d\geq 2$ and most values of $\kappa$. For example, [5] establishes this non triviality for $d=2$ and $\kappa=3$, while also proving that percolation does not occur when $d\geq\kappa=2$, even at $p=1$. Further progress was made in [15], which provides quantitative upper bounds on the critical time and characterizes the phase transition for all $d\geq 3$ and most choices of $\kappa\geq 3$. Nonetheless, it remains open to determine whether the phase transition is trivial ($p_{\mathrm{c}}=1$) for $\kappa\in\{3,\dots,9\}$ for some dimensions, but it is conjectured that this is never the case. In view of this, let us emphasise that Theorem 1 applies regardless of whether the phase transition is trivial.
It is straightforward that the CDP is stochastically dominated by standard Bernoulli percolation for every fixed $p\in[0,1]$. Consequently, Theorem 1 holds trivially whenever $p$ lies below the Bernoulli percolation critical threshold, denoted by $\overline{p_{\mathrm{c}}}$. However, it is known that $p_{\mathrm{c}}>\overline{p_{\mathrm{c}}}$ (see Theorem 1 in [5] for the case $d=2$ and $\kappa=3$; the argument extends for the general case), and therefore Theorem 1 remains nontrivial for all choices of $d$ and $\kappa$.
Let $\widehat{p_{\mathrm{c}}}=\sup\{p\in[0,1]:\sum_{x\in{\mathbb{Z}}^{d}}{\mathbb{P}}(0\xleftrightarrow{p}x)<\infty\}\leq p_{\mathrm{c}}$ denote the suscptibility critical threshold for the CDP. The recent work [6] proves Theorem 1 for $p<\widehat{p_{\mathrm{c}}}$. Our result strengthens their findings: in particular, Theorem 1 implies $p_{\mathrm{c}}=\widehat{p_{\mathrm{c}}}$. This answers [6]*Question 1. in the setting of homogeneous (vertex-independent) constraints $\kappa$.
@@ 1.3 Outline of the proof
A standard approach to proving sharpness in percolation models is the one of [9] using a randomised algorithm (see Section 2.4) and the OSSS inequality [23] (see Section 2.3). Applying this method directly fails. On the one hand, the measure $\omega_{p}\circ{\mathbb{P}}$ is not monotonic and lacks other good properties. On the other hand, if we choose to work with the configuration space $0,1]^{\mathcal{E}}$ instead of $\{0,1\}^{\mathcal{E}}$, edges revealed in the exploration of the occurrence of the one-arm event in (1) do not witness this one-arm event on a smaller scale. For this reason, as in [16], we introduce an alternative one-arm event (see Section 2.1) for which revealment is naturally expressed in terms of its probability. Namely, this event requires a path whose first part (head) is open and whose second part (tail) has decreasing values of the uniform variables on edges. This modified one-arm event satisfies a Russo formula (see Lemma 6).
Putting the above ingredients together, one reaches the main difficulty: the Russo formula features a different pivotality event ($p$-pivotality) from the influences appearing in the OSSS inequality ($U$-pivotality). In Proposition 8, we provide a transfer from one to the other. The statement is largely inspired by the essential enhancement technique of [2]. However, the lack of finite-energy property and the infinite range of dependence make its implementation very delicate (see e.g. [7] for a very different setting where an intricate unbounded range pivotality transfer is also performed). The base idea of the proof of Proposition 8 is simple. The probability that the state of an edge depends on the uniform variables far from it decays super-exponentially, while configuration modifications usually come at an exponential cost. We can then hope to modify the configuration in the neighbourhood of the decreasing cluster of a $U$-pivotal edge in such a way that we produce a $p$-pivotal edge. Unfortunately, there exist clusters whose probability of being decreasing is only exponentially small in their volume (e.g. the fractal ternary square space-filling tree).
@@ 2.2 Transforming pivotality
Write $\tau_{n}(p)={\mathbb{P}}(E_{n}(p))$. The notion of $p$-pivotality is motivated by the following Russo formula (we refer the reader to Proposition 3 in [3] for a related result).
@@
Any good configuration $U^{\prime}$ satisfies $\omega_{p}(U^{\prime},f)=\omega^{+}_{f}$ for all $f\in\mathcal{E}\setminus\mathcal{E}_{e}$, and $\omega_{p}(U^{\prime},f)=\mathbbm{1}_{U^{\prime}_{f}\leq p}$ for $f\in\mathcal{E}_{e}$ (also see Lemma 1 in [3]).
@@
Let $\mathcal{E}_{+}=\{f\in\mathcal{E}_{e}:\omega^{+}_{f}=1\}$. Let $\mathcal{F}$ be a maximal subset of $\mathcal{E}_{+}$, satisfying that no good configuration $U^{\prime}$ with $\omega_{p}(U^{\prime},f)=0$ for $f\in\mathcal{E}_{+}\setminus\mathcal{F}$ belongs to $E_{n}(p)$. By Claims 10 and 11, $\mathcal{F}$ exists and $\mathcal{F}\neq\mathcal{E}_{+}$. Moreover, by the assumption of Case A, no witness of a good configuration in $E_{n}(p)$ can have a tail intersecting $\mathcal{V}_{e}$. Therefore, recalling Claim 9, within good configurations $U^{\prime}$, the event $U^{\prime}\in E_{n}(p)$ depends only on $\mathbbm{1}_{U^{\prime}_{f}\leq p}$ for $f\in\mathcal{E}_{+}$. Therefore, for any good configuration $U^{\prime}$ such that $\mathbbm{1}_{U^{\prime}_{f}\leq p}=\mathbbm{1}_{f\in\mathcal{F}}$ for all $f\in\mathcal{E}_{+}$, we have that each $f\in\mathcal{E}_{+}\setminus\mathcal{F}$ is $p$-pivotal for $U^{\prime}$ for the event $E_{n}(p)$ (recall Definition 5) and, in particular, such $f$ exist.
@@ 2.3 The OSSS inequality
An algorithm $\mathbf{A}$ determining $f$ reveals the coordinates of $x$ sequentially, with each choice depending on the values revealed thus far. The process terminates once the value of $f(x)$ is determined, independently of the unrevealed coordinates. For a formal definition of randomized algorithms, we refer to [23]. Associated to $\mathbf{A}$ and $f$, we define the revealment and influence of coordinate $i\in I$ by
@@
where $x^{i}$ is obtained from $x$ by resampling the $i$-th coordinate independently according to $\pi$, leaving the others unchanged. The OSSS inequality [23] states that for any Boolean function $f:X^{I}\to\{0,1\}$ and any algorithm $\mathbf{A}$ that determines $f$,
@@
This inequality was initially formulated for finite sets $X$ and $I$. Nonetheless, the result remains valid when $X$ is a general space and $I$ is countable (see [8]*Remark 5). Applying (5) to $X=[0,1]$, $\pi$ the Lebesgue measure over $X$ and $I=\mathcal{E}$, so that $\pi^{\otimes I}={\mathbb{P}}$, and the function $f=\mathbbm{1}_{E_{n}(p)}$, and setting $\delta(\mathbf{A})=\max_{e\in\mathcal{E}}\delta_{e}(\mathbf{A})$, we obtain
@@ 2.4 The randomized algorithm approach
In this section, we apply the method of [9]*Section 3 to produce a randomized algorithm with low revealment and deduce a sharpness result for the modified one-arm event $E_{n}(p)$. Since there is little novelty here, the presentation is rather concise and we recommend referring to [9] for more details.
Lemma 14 (Differential inequality implies sharpness [9]*Lemma 3.1).
@@ Acknowledgements
Appendix A Proof of Corollary 3
The proof relies on a one-step renormalisation scheme. For $N,M\in{\mathbb{N}}$ and $x=(x_{1},\dots,x_{d})\in{\mathbb{Z}}^{d}$, define
$B_{N,M}(x)=Nx+[-M,M)^{d},$
the box of side-length $2M$ centered at $Nx$. Fix $x\in{\mathbb{Z}}^{d}$ and consider the events
[…]
We introduce a renormalised lattice, isomorphic to ${\mathbb{Z}}^{d}$, whose vertices correspond to the boxes
$\{B_{N,N}(x),x\in{\mathbb{Z}}^{d}\}.$
Given a configuration $U\in\Omega$ and $x\in{\mathbb{Z}}^{d}$, we say that the box $B_{N,N}(x)$ is good in $U$ if $\mathcal{E}_{N}\cap\mathcal{F}_{N}$ occurs; otherwise it is bad.
We claim that the resulting renormalised process of good and bad boxes is 1-dependent (with respect to the $\|\cdot\|_{\infty}$ norm). To see this, it suffices to show that the event $\mathcal{E}_{N}(x)\cap\mathcal{F}_{N}(x)$ is measurable with respect to $(U_{e})_{e\subset B_{N,N}(x)}$. Indeed, the presence of decreasing paths is clearly measurable and, on $\mathcal{F}_{N}(x)$, the $p$-open edges in $B_{N,\lfloor 3N/4\rfloor}(x)$ only depend on $U_{e}$ for $e\subset B_{N,\lceil N(3/4+1/10)\rceil}\subset B_{N,N}$.
Fixing $x\in{\mathbb{Z}}^{d}$ and $N>10$, let us estimate the probabilities of $\mathcal{E}_{N}(x)$ and $\mathcal{F}_{N}(x)$. By the union bound, translation invariance, and Theorem 1, we obtain
${\mathbb{P}}(\mathcal{E}_{N}^{c})\leq(2N)^{d}e^{-\alpha_{p}N/10}.$ (7)
Similarly, by the union bound, we have
${\mathbb{P}}(\mathcal{F}_{N}^{c})\leq\frac{(2N)^{d}2d(2d-1)^{\lfloor N/10\rfloor-1}}{\lfloor N/10\rfloor!}.$ (8)
Combining (7) and (8), we conclude that
$\lim_{N\to\infty}{\mathbb{P}}(B_{N,N}(x)\mbox{ is good})=1.$
By the Liggett-Schonmann-Stacey theorem [21], one can choose $N$ sufficiently large such that this 1-dependent renormalised process stochastically dominates a supercritical independent site percolation model. For such a model, the size of the cluster of vertices corresponding to bad boxes (which is subcritical) is known to have an exponential tail (see Theorem 6.75 in [14]). The result then follows from the observation that the $p$-open cluster $C_{0}$ of the origin is contained in the box $B_{N,N}(0)$ together with the cluster of bad boxes containing the origin of the renormalised lattice.